OpenStudy (anonymous):

Use the unit circle, along with the definitions of the circular functions, to find the exact value for the function, s= 8(pi)/3, what does secant 8(pi)/3 equal?

OpenStudy (jdoe0001):

well, use your Unit Circle to get the x,y values for the angle \(\bf \cfrac{8\pi}{3}\) keep in mind that \(\bf \cfrac{8\pi}{3}\implies 2\cfrac{2\pi}{3}\implies 2\pi + \cfrac{2\pi}{3}\) and also that \(\bf sec(\theta)=\cfrac{1}{cos(\theta)}\)

OpenStudy (anonymous):

Thank you! :)